Assume that you plan to use a significance level of α = 0.05 to test the claim that p1 = p2. Use the given sample sizes and numbers of successes to find the pooled estimate. Round your answer to the nearest thousandth.

n1 = 677 n2 = 3307
x1 = 172 x2 = 564

Group of answer choices

a) 0.223

b) 0.118

c) 0.185

d) 0.204
Answer #1

The pooled estimate is ,

\hat{P}=\frac{X_1+X_2}{n_1+n_2}=\frac{172+564}{677+3307}=0.185